\begin{thebibliography}{10}

%\bibitem[AM]{AM}
%A. Adem and R. J. Milgram,
%``Cohomology of Finite Groups",
%Grundlehren der mathematischen Wissenschaften Vol 309,
%Springer-Verlag, Berlin, 1994.

%\bibitem[ASZ]{ASZ}
%K. Ajitabh, S.P. Smith and J.J. Zhang,
%Auslander-Gorenstein rings, Comm. Algebra
%{\bf 26} (1998), no. 7, 2159--2180.

%\bibitem[AD]{AD} J. Alev and F. Dumas,  Op\'{e}rateurs diff\'{e}rentiels invariants et probl\`{e}me de Noether. (French) [Invariant differential operators and the Noether problem] Studies in Lie theory, 21–50, Progr. Math., 243, Birkhäuser Boston, Boston, MA, 2006.
\bibitem[AP]{AP}
J. Alev and P. Polo,
A rigidity theorem for finite group actions on enveloping algebras of
semisimple Lie algebras,
Adv. Math. {\bf 111} (1995), no. 2, 208--226.

%\bibitem[Al]{Al}
%A. Aljouiee,
%On weak crossed products, Frobenius algebras, and the weak Bruhat ordering,
%J. Algebra {\bf 287}  (2005),  no. 1, 88--102.

%\bibitem[AP]{AP}
%J. Alev and P. Polo, A rigidity theorem for finite group actions
%on enveloping algebras of semisimple Lie algebras, Adv. Math. {\bf
%111} (1995), no. 2, 208--226.

\bibitem[AS]{AS}
 M. Artin and W. F. Schelter,
Graded algebras of global dimension $3$,
Adv. in Math. {\bf 66} (1987), no. 2,
171--216.

%\bibitem[ASt1]{ASt1}
%M. Artin and J. T. Stafford,  Noncommutative graded
%           domains with quadratic growth, Invent. Math. {\bf 122}
%           (1995), no. 2, 231--276.

%\bibitem[ASt2]{ASt2}
%M. Artin and J. T. Stafford, Semiprime graded algebras
%           of dimension two, preprint.

%\bibitem[ATV]{ATV}%ATV1
%M. Artin, J. Tate and M. Van den Bergh,
%Some algebras associated to automorphisms of elliptic
%curves, ``The Grothendieck Festschrift,''
%Vol. I, ed.\ P. Cartier et al., Birkh\"{a}user Boston
%1990, 33-85.

%\bibitem{[ATV2]} M. Artin, J. Tate and M. Van den Bergh,
%         Modules over regular algebras of dimension $3$, Invent. Math.
%         {\bf 106} (1991), no. 2, 335--388.

%\bibitem[AZ]{AZ}
%M. Artin and J. J. Zhang, Noncommutative projective
%schemes, Adv. Math. {\bf 109} (1994), 228-287.

%\bibitem[BM]{BM}
%G. Barou and M.-P. Malliavin,
%Sur la r{\'e}solution injective minimale de l'alg{\`e}bre enveloppante d'une
%alg{\`e}bre de Lie r{\'e}soluble, (French)
%J. Pure Appl. Algebra 37 (1985), no. 1, 1--25.

%\bibitem[BR]{BR} G. Benkart and T. Roby, Down-up algebras, {\it J.
%Algebra} 209 (1998), 305-344. Addendum,  {\it J. Algebra} {\bf
%213}  (1999), no. 1, 378.

\bibitem[BB]{BB} T. Banica and J. Bichon, 
Hopf images and inner faithful representations, Glasg. Math.
J. {\bf 52} (2010), no. 3, 677-703.

\bibitem[BHZ]{BHZ}
Y.-H. Bao, J.-W. He and J.J. Zhang,
\emph{Pertinency of Hopf actions and quotient categories
of Cohen-Macaulay algebras}, (2016) preprint, arXiv:1603.02346

%\bibitem[BHZ2]{BHZ2}
%Y.-H. Bao, J.-W. He and J.J. Zhang,
%\emph{Noncommutative Auslander theorem},
%in preparation (2016).


%\bibitem[Ba]{Ba}
%Y. Bazlov,
%Nichols-Woronowicz algebra model for Schubert calculus on Coxeter groups,
%J. Algebra {\bf 297}  (2006),  no. 2, 372--399.



%\bibitem[BB]{BB}
%Y. Bazlov and A. Berenstein,
%Mystic reflection groups,
%SIGMA Symmetry Integrability Geom. Methods Appl. {\bf 10}
%(2014), Paper 040, 11 pp.

%\bibitem[Be]{Be}
%D.J. Benson, Polynomial invariants of finite groups, London
%Mathematical Society Lecture Note Series, 190. Cambridge
%University Press, Cambridge, 1993.

\bibitem[BR]{BR}
G. Benkart and T. Roby,
Down-up algebras,
J. Algebra {\bf 209} (1998), no. 1, 305-344.

\bibitem[Be]{Be}
G.M. Bergman,
The diamond lemma for ring theory,
Adv. in Math. {\bf 29} (1978), no. 2, 178-218.

%\bibitem[BjB]{BjB}
%A. Bj\"{o}rner and F. Brenti,
%Combinatorics of Coxeter groups,
%Graduate Texts in Mathematics, {\bf 231}, Springer, New York, 2005.

%\bibitem[Bo1]{Bo1}
%R. Bocklandt,
%Graded Calabi Yau algebras of dimension 3,
%J. Pure Appl. Algebra {\bf 212} (2008), no. 1, 14--32.

%\bibitem[BSW]{BSW}
%R. Bocklandt, T. Schedler and M. Wemyss,
%Superpotentials and higher order derivations,
%J. Pure Appl. Algebra {\bf 214} (2010), no. 9, 1501--1522.

%\bibitem[BRR]{BRR}	
%F. Brenti, V. Reiner and and Y. Roichman,
%Alternating subgroups of Coxeter groups,
%J. Combin. Theory Ser. A {\bf 115} (2008), no. 5, 845--877.

%\bibitem[BW]{BW}
%F. Brenti and V. Welker,
%The Veronese construction for formal power series and graded algebras,
%Adv. in Appl. Math., {\bf 42}, no. 4, 545-556.

%\bibitem[BZ]{BZ}
%K.A. Brown and J.J. Zhang,
%Dualising complexes and twisted Hochschild (co)homology for
%Noetherian Hopf algebras,
%J. Algebra, {\bf 320} (2008), no. 5, 1814--1850.

%\bibitem[CA]{CA}
%J. Cameron Atkins, The invariant theory of $k_{-1}[x_1,\cdots,x_n]$
%under permutation reprensentations, Mater Thesis, Wake Forest
%Unibersity, 2012.

%\bibitem[CF]{CF} J. W. Cassels and A. Fr\"{o}hlich, Algebraic Number Theory
%in {\it Proceedings of an Instructional Conference}, Academic Press, New York,
%1967.

%\bibitem[CE]{CE}
%A. Clark and J. Ewing, The realization of polynomial algebras as
%cohomology rings, Pacific J. Math. {\bf 50} (1974), 425--434.

%\bibitem[Co]{Co}
%H.S.M. Coxeter, Discrete groups generated by reflections, Ann. of
%Math. (2) {\bf 35} (1934), no. 3, 588--621.

%\bibitem[CS]{CS}
%T. Cassidy and B Sheldon, Generalizing the notion of Koszul algebra,
%Math. Z. 260 (2008) 93-114.

%\bibitem[CV]{CV}
%T. Cassidy and M. Vancliff, Generalizations of graded Clifford
%algebras and of complete intersections, 2008 preprint.

%\bibitem[CWZ]{CWZ}
%K. Chan, C. Walton and J.J. Zhang,
%Hopf actions and Nakayama automorphisms,
%J. Algebra {\bf 409}  (2014), 26--53.


%\bibitem[CKWZ]{CKWZ}
%K. Chan, E. Kirkman, C. Walton and J.J. Zhang,
%Quantum binary polyhedral groups
%and their actions on quantum planes, ArXiv 1303.7203,
%to appear in J. Reine Angew. Math.
%(Crelle’s Journal).

\bibitem[CKWZ1]{CKWZ1}
K. Chan, E. Kirkman, C. Walton and J.J. Zhang, 
McKay correspondence for semisimple Hopf actions 
on regular graded algebras I, in preparation (2016).

\bibitem[CKWZ2]{CKWZ2}
K. Chan, E. Kirkman, C. Walton and J.J. Zhang, 
McKay correspondence for semisimple Hopf actions 
on regular graded algebras II, in preparation  (2016).

%\bibitem[Ch]{Ch}
%C. Chevalley,
%Invariants of finite groups generated by reflections, Amer. J. Math
%{\bf 77} (1955), 778--782.

%\bibitem[CM]{CM}
%M. Cohen and S. Montgomery,
%Group-graded rings, smash products, and group actions,
%Trans. Amer. Math. Soc. {\bf 282}  (1984),  no. 1, 237--258.


%\bibitem[DK]{DK}
%H. Derksen, G. Kemper, ``Computational Invariant Theory",
%Encyclopedia of Mathematical Sciences 130: Invariant Theoy
%and Algebraic Transformation Groups I,
%Springer-Verlag, Berlin, 2002.

%\bibitem[EW]{EW}
%P. Etingof and C. Walton,
%Semisimple Hopf actions on commutative domains,
%Adv. Math {\bf 251} (2014), 47--61.



%\bibitem[DH]{DH} M. Domokos and P. Heged\"{u}s,
%Noether's bound for polynomial invariants of
%finite groups, Arch. Math. 74 (2000), 161-167.

%\bibitem[EFOS]{EFOS} F. Eshmatov, V. Futorny, S. Ovsienko, and J.F. Schwarz, Noncommutative Noether's problem for complex reflection groups, (2015 preprint) arXiv:1505.05626.
%\bibitem[FHT]{FHT}
%Y. F{\' e}lix, S. Halperin and J.-C. Thomas,
%Elliptic Hopf algebras,
%J. London Math. Soc. (2) {\bf 43} (1991), no. 3, 545--555.

%\bibitem[FS]{FS}
%S. Fomin and R.P. Stanley,
%Schubert Polynomials and the Nilcoxeter Algebra,
%Adv. Math {\bf 103} (1994), 196--207.

%\bibitem[GR]{GR}
%J. Gaddis and D. Rogalski,
%in preparation.

%\bibitem[GN]{GN}
%C. Galindo and S. Natale,
%Simple Hopf algebras and deformations of finite groups,
%Math. Res. Lett. {\bf 14}  (2007),  no. 6, 943--954.

%\bibitem[Go]{Go}
%M. G\"{o}bel,
%Computing bases for rings of permutation-invariant polynomials,
%J. Symbolic Comput. {\bf 19} (1995), 285-291.

%\bibitem[G1]{G1}
%N.L. Gordeev,
%Invariants of linear groups generated by matrices with two
%nonidentity eigenvalues,
%Zap. Nauchn. Sem. Leningrad. Otdel. Mat. Inst. Steklov.
%(LOMI)  {\bf 114} (1982),
%120-130; English translation in J. Soviet Math. {\bf 27} (1984), no. 4.

%\bibitem[G2]{G2}
%N.L. Gordeev,
%Finite linear groups whose algebra of invariants is a complete intersection
%(Russian), Izv. Akad. Nauk SSR Ser. Mat. {\bf 50} (1986), no. 2, 343-392.
%(English translation: Math. USSR-Inv. {\bf 28}(1987), no. 2, 335-379.)

%\bibitem[Gu]{Gu}
% T.H. Gulliksen,
%A homological characterization of local complete intersections,
%Compositio Math. 23 (1971), 251-255.

%\bibitem[He]{He}
%H. Helfgott, Growth in groups: ideas and perspectives,
%Bull. Amer. Math. Soc. (N.S.) {\bf 52} (2015), no. 3, 357--413

%\bibitem[HW]{HW} G. H. Hardy and E. M Wright, ``An Introduction to the
%Theory of Numbers", Fourth Edition, Oxford University Press,
%London, 1960.

%\bibitem[JiZ]{JiZ}
%N. Jing and J.J. Zhang, On the trace of graded automorphisms, J.
%Algebra {\bf 189} (1997), no. 2, 353--376.

%\bibitem[JiZ2]{JiZ2}
%N. Jing and J.J. Zhang, Gorensteinness of invariant subrings of
%quantum algebras, J. Algebra {\bf 221} (1999), no. 2, 669--691.

%\bibitem[{\bf JW}]{JW} D. A. Jordan and I.E. Wells, Invariants for
%automorphisms of certain iterated skew polynomial
%rings, {\it Proc. Edinburgh Math. Soc.} {\bf 39} (1996), 461-472.

%\bibitem[J1]{J1}
%P. J{\o}rgensen,
%Non-commutative graded homological identities,
%J. London Math. Soc. (2) {\bf 57}  (1998),  no. 2, 336--350.

%\bibitem[J2]{J2}
%P. J{\o}rgensen,
%Local cohomology for non-commutative graded algebras,
%Comm. Algebra {\bf 25}  (1997),  no. 2, 575--591.


%\bibitem[JoZ]{JoZ}
%P. J{\o}rgensen and J.J. Zhang, Gourmet's guide to Gorensteinness,
%Adv. Math. {\bf 151} (2000), no. 2, 313--345.

%\bibitem[Jo]{Jo}
%A. Joseph, Coxeter structure and finite group action, Alg{\`e}bre
%non commutative, groupes quantiques et invariants (Reims, 1995),
%185--219, S{\'e}min. Congr., 2, Soc. Math. France, Paris, 1997.

%\bibitem[GW]{GW}
%K.R. Goodearl and  R.B. Warfield Jr.,
%``An introduction to noncommutative Noetherian rings'',
%Second edition. London Mathematical Society Student Texts, 61.
%Cambridge University Press, Cambridge, 2004.

%\bibitem[KW]{KW} V. Kac and K. Watanabe, Finite linear groups whose ring
%of invariants is a complete intersection, Bull, Amer. Math. Soc. (N.S.) {\bf 6}
%(1982), no. 2, 221-223.

%\bibitem[KM]{KM}
%A.N. Kirillov and T. Maeno,
%Nichols-Woronowicz model of coinvariant algebra of complex reflection groups,
%J. Pure Appl. Algebra  {\bf 214}  (2010),  no. 4, 402--409.

\bibitem[Ki]{Ki}
E. Kirkman, Invariant theory of Artin-Schelter regular algebras:
a survey, preprint, (2015), arXiv:1506.06121v1.

\bibitem[KK]{KK}
E. Kirkman and J. Kuzmanovich,
Fixed subrings of Noetherian graded regular rings,
J. Algebra {\bf 288} (2005), no. 2, 463--484.


\bibitem[KKZ1]{KKZ1} E. Kirkman, J. Kuzmanovich, and J.J. Zhang,
Rigidity of graded regular algebras,
Trans. Amer. Math. Soc. {\bf 360} (2008), no. 12, 6331--6369.

\bibitem[KKZ2]{KKZ2}
E. Kirkman, J. Kuzmanovich and J.J. Zhang, A
Shephard-Todd-Chevalley theorem for noncommutative regular
algebras, Algebr. Represent. Theory {\bf 13} (2010), no. 2, 127--158

\bibitem[KKZ3]{KKZ3}
E. Kirkman, J. Kuzmanovich and J.J. Zhang, Gorenstein subrings of invariants
under Hopf algebra actions, J. Algebra {\bf 322} (2009), no. 10, 3640--3669.

%\bibitem[KKZ4]{KKZ4}
%E.Kirkman, J. Kuzmanovich, and J.J. Zhang,
%Noncommutative complete intersections,
%J. Algebra {\bf 429} (2015), 253--286.

\bibitem[KKZ4]{KKZ4}
E.Kirkman, J. Kuzmanovich, and J.J. Zhang,
Invariant theory of finite group actions on down-up algebras,
Transform. Groups {\bf 20} (2015), no. 1, 113--165.

\bibitem[KKZ5]{KKZ5}
E.Kirkman, J. Kuzmanovich, and J.J. Zhang,
Nakayama automorphism and rigidity of dual reflection group
coactions, (2016) arXiv:1605.01958.


%\bibitem[KKZ7]{KKZ7}
%E.Kirkman, J. Kuzmanovich, and J.J. Zhang,
%Dual reflection groups: Examples, in preparation.

\bibitem[KMP]{KMP} E. Kirkman, I. Musson, D. Passman, Noetherian
down-up algebras, {\it Proc. Amer. Math. Soc.} {\bf 127} (1999),
3161-3167.

%\bibitem[KL]{KL}
%G. Krause and T. Lenagan, ``Growth of Algebras and
%Gelfand-Kirillov Dimension, revised edition, Graduate Studeis in
%Mathematics, Vol. 22, AMS, Providence, 2000.

%\bibitem[Le]{Le}
%T. Levasseur,
%Some properties of noncommutative regular rings,
%Glasgow Math. J. {\bf 34} (1992), 277-300.

%\bibitem[LMZ]{LMZ}
%J.-F. L\"u, X.-F. Mao and J.J. Zhang,
%Nakayama automorphism and applications,
%Trans. Amer. Math. Soc., (to appear), preprint
%arXiv:1408.5761.

%\bibitem[LMZ2]{LMZ2}
%J.-F. L\"u, X.-F. Mao and J.J. Zhang,
%Nakayama automorphism of a class of graded algebras,
%\emph{Israel Journal of Mathematics}, (to appear), preprint
%arXiv:1512.00581.


%\bibitem[MR]{MR}
%J. C. McConnell and J. C . Robson, ``Noncommutative Noetherian
%Rings,'' Wiley, Chichester, 1987.

%\bibitem[MS]{MS}
%A. Milinski and H.-J. Schneider,
%Pointed indecomposable Hopf algebras over Coxeter groups,
%New trends in Hopf algebra theory (La Falda, 1999),  215–236,
%Contemp. Math., {\bf 267}, Amer. Math. Soc., Providence, RI, 2000.


%\bibitem[Mon]{Mon}
%S. Montgomery,
%``Hopf algebras and their actions on rings'',
%         CBMS Regional Conference Series in
%Mathematics, {\bf 82},
 %        Providence, RI, 1993.
%\bibitem[Mo]{Mo} R. Mollin, ``Algebraic Number Theory".
%Chapman \& Hall/CRC, Boca Raton, FL, 1999.

%\bibitem[Mon]{Mon}
%S. Montgomery, ``Fixed rings of finite automorphism groups of
%associative rings'', Lecture Notes in Mathematics, 818. Springer,
%Berlin, 1980

%\bibitem[N1]{N1} H. Nakajima, Rings of invariants of finite groups which
%are hypersurfaces, J. Algebra {\bf 80} (1983), 279-294.

%\bibitem[N2]{N2} H. Nakajima, Rings of invariants of finite groups which
%are hypersurfaces II, Adv. Math.

%\bibitem[N3]{N3}
%H. Nakajima, Quotient singularities which are complete intersections,
%Manuscripta Math. {\bf 48} (1984), no.1-3, 163-187.

%\bibitem[N4]{N4} H. Nakajima,
%Quotient complete intersections of affine spaces by finite linear groups,
%Nagoya Math J. {\bf 98} (1985), 1-36.

%\bibitem[NW]{NW}
%H. Nakajima and K. Watanabe,
%``Complete intersections" (Acireale, 1983), 102-120,
%Lecture Notes in Math. 1092, Springer, Berlin, 1984.

%\bibitem[Ne]{Ne} M. Neusel, ``Invariant Theory",
%Student Mathematical Library Vol. 36, American
%Mathematical Society, Providence, R.I, 2007.

%\bibitem[Ni]{Ni}
%D. Nikshych,
%$K_0$-rings and twisting of finite dimensional semisimple Hopf algebras,
%Commun. Algebra {\bf 26} (1998), 321--342; Commun. Algebra {\bf 26} (1998),
%2019.

%\bibitem[No]{No}
%E. Noether, Der Endlichkeitssatz der Invarianten endlicher Gruppen,
%Math. Ann. 77 (1916), 89-92.

%\bibitem[NV]{NV}
%C. N\u ast\u asescu, F. van Oystaeyen,
%``Graded ring theory'',
%         North-Holland Mathematical Library, 28.
% North-Holland Publishing
%         Co., Amsterdam-New York, 1982.

%\bibitem[Pr]{Pr}
%S.B. Priddy,
%Koszul resolutions,
%Trans. Amer. Math. Soc. {\bf 152} (1970) 39--60.


%\bibitem[RRZ1]{RRZ1}
%Z. Reichstein, D. Rogalski and J.J. Zhang,
%Projectively simple rings,
%Adv. in Math, {\bf 203} (2006), 365--407.

%\bibitem[RRZ2]{RRZ2}
%M. Reyes, D. Rogalski and J.J. Zhang,
%Skew Calabi-Yau algebras and homological identities,
%Adv. Math. {\bf 264} (2014), 308--354.

%\bibitem[RRZ3]{RRZ3}
%M. Reyes, D. Rogalski and J.J. Zhang,
%Skew Calabi-Yau triangulated categories and Frobenius Ext-algebras,
%Trans. Amer. Math. Soc., (to appear), preprint (2014)
%arXiv:1408.0536.


%\bibitem[Ro]{Ro}
%A. Rotbart,
%Generator sets for the alternating group,
%S{\'e}m. Lothar. Combin. {\bf 65} (2010/12), Art. B65b, 16 pp.


%\bibitem[S]{S}
%B.J. Schmid, Finite groups and invariant theory, in;
%P. Dubreil, M.-P. Malliavin, eds.,
%Topics in Invariant Theory, Lecture Notes Math. 1478, Springer-Verlag,
%Berlin, 1991.

\bibitem[ST]{ST}
G.C. Shephard and J.A. Todd, Finite unitary reflection groups,
Canadian J. Math. {\bf 6}, (1954). 274--304.


%\bibitem[SW]{SW}
%L. W. Small and R. B. Warfield Jr.,
%Prime affine algebras
%        of Gelfand-Kirillov dimension one. J. Algebra
% {\bf 91} (1984),
%        no. 2, 386--389.


%\bibitem[SSW]{SSW}
%L. W. Small, J.T. Stafford and R. B. Warfield Jr.,
%Affine algebras of Gelfand-Kirillov dimension one are
%PI,
%Math. Proc. Cambridge Philos. Soc. {\bf 97} (1985),
%no. 3, 407--414.

%\bibitem[Sm1]{Sm1}
%S.P. Smith, Can the Weyl algebra be a fixed ring? Proc. Amer.
%Math. Soc. {\bf 107} (1989), no. 3, 587--589.

%\bibitem[Sm1]{Sm1}
%S.P. Smith, Some finite-dimensional algebras related to elliptic
%curves, Representation theory of algebras and related topics
%(Mexico City, 1994), 315--348, CMS Conf. Proc., 19, AMS,
%Providence, RI, 1996.

%\bibitem[SS]{SS}
%S. P. Smith and J. T. Stafford,
%Regularity of the
%four-dimensional Sklyanin algebra,
%Compositio Math. {\bf 83} (1992), no. 3, 259--289.

%\bibitem[SZ]{SZ}
%J. T. Stafford and J. J. Zhang,
%Homological properties of (graded) Noetherian PI
%rings,
%J. Algebra {\bf 168} (1994), no. 3, 988--1026.

%\bibitem[St1]{St1}
%R.P. Stanley,
%Relative invariants of finite
%groups generated by pseudo-reflections, J. Algebra {\bf 49} (1977),
%134-148.

%\bibitem[S1]{S1}
%R.P. Stanley,
%Invariants of finite groups and their applications
%to combinatorics, Bulletin of the AMS, {\bf 1} (1979), 475-511.

%\bibitem[S2]{S2}
%R.P. Stanley,
%Hilbert functions of graded algebras,
%Adv. Math. {\bf 28} (1978), 57-83.



%\bibitem[SV]{SV}
%D. Stefan and C. Vay,
%The cohomology ring of the 12-dimensional Fomin-Kirillov algebra,
%Adv. Math. {\bf 291} (2016), 584–620.

\bibitem[Ste]{Ste}
D.R. Stephenson,
Artin-Schelter regular algebras of global dimension three,
J. Algebra {\bf 183} (1996), no. 1, 55--73.

%\bibitem[StZ]{StZ}
%D.R. Stephenson and J.J. Zhang, Growth of graded Noetherian rings,
%Proc. Amer. Math. Soc. {\bf 125} (1997), no. 6, 1593--1605.

%\bibitem[St]{St} B. Sturmfels, ``Algorithms in Invariant Theory", Texts and Monographs
%in Symbolic Computation, Springer-Verlag , New York, 1993.

%\bibitem[TV]{TV}
%J. Tate and M. Van den Bergh,
%Homological properties of Sklyanin algebras,
%Invent. Math. {\bf 124} (1996), no. 1-3, 619--647.

%\bibitem[VdB1]{VdB1}
%M. Van den Bergh,
%Existence theorems for dualizing complexes over noncommutative
%graded and filtered rings, J. Algebra {\bf 195} (1997), 662--679.

%\bibitem[VdB2]{VdB2}
%M. Van den Bergh,
%Calabi-Yau algebras and superpotentials,
%Selecta Math. (N.S.) {\bf 21} (2015), no. 2, 555--603.

%\bibitem[VdB3]{VdB3}
%M. Van den Bergh,
%Non-commutative crepant resolutions,
%The legacy of Niels Henrik Abel, 749-770, Springer, Berlin,
%{\bf 2004}.

%\bibitem[W1]{W1}
%K. Watanabe, Certain invariant subrings are Gorenstein, I, Osaka J. Math. 11
%(1974), 1-8.

%\bibitem[W2]{W2}
%K. Watanabe, Certain invariant subrings are Gorenstein, II, Osaka J. Math. 11
%(1974), 379-388.

%\bibitem[WR]{WR} K. Watanabe and D. Rotillon, Invariant
%subrings of $\mathbb{C}[x,y,z]$ which are complete
%intersections, Manuscripta Math. {\bf 39} (1982),
%339-357.

%\bibitem[W3]{W3}
%K. Watanabe, Invariant subrings which are complete intersections I
%(Invariant subrings of finite abelian groups),
%Nagoya Math. J. {\bf 77} (1980), 89-98.

%\bibitem[WZ]{WZ}
%Q.-S. Wu and J.J. Zhang,
%Noetherian PI Hopf algebras are Gorenstein,
%Trans. Amer. Math. Soc. {\bf 355} (2003), no. 3, 1043--1066


%\bibitem[WZ4]{WZ4}
%Q.-S. Wu and J.J. Zhang,
%Dualizing complexes over noncommutative local rings,
%J. Algebra, {\bf 239} (2001), 513-548.


%\bibitem[Ye1]{Ye1}
%A. Yekutieli,
%         Dualizing complexes over noncommutative
%graded algebras,
%         J. Algebra {\bf 153} (1992), 41-84.

%\bibitem[Ye2]{Ye2}
%A. Yekutieli,
%The residue complex of a
%         noncommutative graded algebra, J. Algebra
%{\bf 186}
 %        (1996), 522-543.

%\bibitem[YZ1]{YZ1}
%A. Yekutieli and J.J. Zhang,
%Residual complex over
%noncommutative rings, J. Algebra {\bf 259}
%(2003), no. 2,
%451--493.

%\bibitem[YZ]{YZ}
%A. Yekutieli and J.J. Zhang,
%Rings with Auslander dualizing complexes,
%J. Algebra {\bf 213} (1999), no. 1, 1--51.


\end{thebibliography}

\documentclass{amsart}
\usepackage{latexsym,amsxtra,amscd,ifthen}
\usepackage{amsfonts}
\usepackage{verbatim}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amssymb}
\numberwithin{equation}{section}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{conjecture}[theorem]{Conjecture}

\theoremstyle{definition}

\newtheorem{example}[theorem]{Example}

\makeatletter              
\let\c@equation\c@theorem  
\makeatother

\newcommand{\DD}{\mathbb{D}}
\newcommand{\FF}{\mathbb{F}}
\newcommand{\HH}{\mathbb{H}}

\newcommand{\be}{\begin{enumerate}}
\newcommand{\ee}{\end{enumerate}}
\newcommand{\bq}{\begin{eqnarray*}}
\newcommand{\eq}{\end{eqnarray*}}
\newcommand{\bqn}{\begin{eqnarray}}
\newcommand{\eqn}{\end{eqnarray}}

\begin{document}
\title[Rigidity of down-up algebras]
{Rigidity of down-up algebras \\
with respect to finite group coactions}

\author{J. Chen, E. Kirkman and J.J. Zhang}

\address{Chen: School of Mathematical Science,
Xiamen University, Xiamen 361005, Fujian, China}

\email{chenjianmin@xmu.edu.cn}

\address{Kirkman: Department of Mathematics,
P. O. Box 7388, Wake Forest University, Winston-Salem, NC 27109}

\email{kirkman@wfu.edu}

\address{Zhang: Department of Mathematics, Box 354350,
University of Washington, Seattle, Washington 98195, USA}

\email{zhang@math.washington.edu}

\begin{abstract}
If $G$ is a nontrivial finite group coacting on a
graded noetherian down-up algebra $A$ inner
faithfully and homogeneously, then the fixed subring
$A^{co\; G}$ is not isomorphic to $A$. Therefore
graded noetherian down-up algebras are rigid with
respect to finite group coactions, in the sense of
Alev-Polo. An example is given to show that this 
rigidity under group coactions does not have all 
the same consequences as the rigidity under group 
actions.
\end{abstract}

\subjclass[2010]{16E10, 16W22}




\keywords{Down-up algebra,
coaction, rigidity, Artin-Schelter regular algebra, homological determinant} 


\maketitle


\setcounter{section}{-1}

\section{Introduction}
\label{xxsec0}

Throughout this paper, let $\Bbbk$ be a base field that
is algebraically closed of characteristic zero, and let all vector
spaces, (co)algebras, and morphisms be over $\Bbbk$.  

A remarkable theorem of Alev-Polo \cite[Theorem 1]{AP} states:

{\it Let ${\mathfrak g}$ and ${\mathfrak g}'$
be two semisimple Lie algebras. Let $G$ be a finite group of algebra
automorphisms of the universal enveloping algebra $U({\mathfrak g})$
such that the fixed subring $U({\mathfrak g})^G$ is isomorphic to 
$U({\mathfrak g}')$. Then $G$ is trivial and ${\mathfrak g}\cong 
{\mathfrak g}'$.}

Alev-Polo called this result a rigidity theorem for universal 
enveloping algebras. In addition, they proved a rigidity theorem for 
the Weyl algebras \cite[Theorem 2]{AP}. Kuzmanovich and the 
second- and third-named authors proved Alev-Polo's rigidity theorems 
in the graded case in \cite[Theorem 0.2 and Corollary 0.4]{KKZ1}.

 (Commutative) polynomial rings are not rigid; indeed,
by the classical Shephard-Todd-Chevalley Theorem if $G$ is a 
reflection group acting on a commutative polynomial ring $A$ 
then $A^G$ is isomorphic to $A$. Artin-Schelter
regular algebras \cite{AS} are considered to be a natural analogue of
polynomial rings in many respects. This paper concerns a
class of noncommutative Artin-Schelter regular algebras. The 
rigidity of a noncommutative algebra is closely related
to the lack of reflections in the noncommutative setting \cite{KKZ1}.
Therefore the rigidity of an algebra leads to a trivialization of the
Shephard-Todd-Chevalley theorem \cite{ST, KKZ2}, which is one of the key
results in noncommutative invariant theory \cite{Ki}.  The rigidity property is
also related to Watanabe's criterion for the Gorenstein property, see
\cite[Theorem 4.10]{KKZ3}. Some recent work in noncommutative algebraic
geometry connects the rigidity property and the lack of reflections to Auslander's
theorem \cite{BHZ}, which is one of the fundamental ingredients in the 
McKay correspondence \cite{CKWZ1, CKWZ2}. Further understanding of the rigidity 
property will have implications for several other research directions.

In \cite{KKZ5}, rigidity with respect to group coactions is
studied. Let $A$ be a connected ($\mathbb{N}$-)graded $\Bbbk$-algebra. 
A $G$-coaction on $A$ (preserving the ${\mathbb N}$-grading) is 
equivalent to a $G$-grading of $A$ (compatible with the original
${\mathbb N}$-grading), and the fixed subring $A^{co G}$ is $A_e$, 
the component of the unit element $e \in A$ under the $G$-grading. We 
recall a definition \cite[Definition 0.8]{KKZ5}:
 we say that a connected graded algebra $A$ is {\it rigid with
respect to group coactions} if for every nontrivial finite group $G$
coacting on $A$ homogeneously and inner faithfully, the fixed subring
$A^{co\; G}$ is not isomorphic to $A$ as algebras. The following
Artin-Schelter regular algebras are rigid with respect to group coactions
\cite[Theorem 0.9]{KKZ5}:
\begin{enumerate}
\item[(a)]
The homogenization of the universal enveloping algebra of a finite dimensional
semisimple Lie algebra.
\item[(b)]
The Rees ring of the Weyl algebra with respect to the standard
filtration.
\item[(c)]
The non-PI Sklyanin algebras of global dimension at least 3.
\end{enumerate}

These results can be viewed as dual versions of the rigidity theorems
proved in \cite[Theorem 0.2 and Corollary 0.4]{KKZ1}.

Down-up algebras were introduced by Benkart-Roby in \cite{BR}
as a tool to study the structure of certain posets. Graded
noetherian down-up algebras are Artin-Schelter regular algebras of
global dimension three with two generators \cite{KMP}. We 
recall the definition of only a graded noetherian down-up algebra.  For
$\alpha$ and $\beta$ scalars in $\Bbbk$, let the {\it down-up
algebra} $\DD(\alpha,\beta)$ be the algebra
generated by $u$ and $d$ and subject to the relations 
\begin{align}
\label{E0.0.1}\tag{E0.0.1}
u^2 d &=\alpha udu +\beta d u^2,\\
\label{E0.0.2}\tag{E0.0.2}
u d^2 &=\alpha dud +\beta d^2 u.
\end{align}
When $\alpha=0$ we denote the down-up algebra $\DD(0, \beta)$ by 
$\DD_{\beta}$. In this
paper we always assume that $\beta\neq 0$, or equivalently,
$\DD(\alpha,\beta)$ is a graded noetherian Artin-Schelter
regular algebra of global dimension three. The groups of
algebra automorphisms of down-up algebras (which depend upon the values 
of $\alpha$ and $\beta$) were computed in
\cite{KK}. These groups are rich enough to provide many nontrivial
examples. Some invariant theoretic aspects of down-up
algebras have been studied in \cite{KK, KKZ4}.

There is a rigidity result concerning group actions on down-up
algebras, see \cite[Proposition 6.4]{KKZ1}. The only theorem in this
paper is the following rigidity result for group coactions on graded
noetherian down-up algebras.

\begin{theorem}
\label{xxthm0.1} Let $A$ be a graded noetherian down-up algebra
$\DD(\alpha,\beta)$ and let $G$ be a nontrivial finite group
coacting on $A$ inner faithfully and homogeneously. Then the
fixed subring $A^{co\; G}$ is not Artin-Schelter regular.
As a consequence, $A$ is rigid with respect to finite group
coactions.
\end{theorem}

One can also consider general Hopf algebra actions on
the down-up algebras. The following conjecture is reasonable.

\begin{conjecture}
\label{xxcon0.2}
The graded noetherian down-up algebras are rigid with respect
to semisimple Hopf algebra actions in the sense of
\cite[Remark 0.10]{KKZ5}.
\end{conjecture}

In a slightly different language, Theorem \ref{xxthm0.1} says 
that graded noetherian down-up algebras do not admit 
``dual reflection groups'' for group coactions in the sense of
\cite[Definition 0.1]{KKZ5}. A result in \cite[Corollary 4.11]{KKZ3}
states: {\it Let $B$ be a noetherian Artin-Schelter 
regular domain, and let $G$ be a finite group acting on $B$
homogeneously. Suppose that $G$ contains no quasi-reflection. 
Then the fixed subring $B^G$ is Artin-Schelter Gorenstein 
if and only if the homological determinant of the $G$-action
is trivial.} When $A$ is a graded noetherian down-up algebra,
$A$ does not have a quasi-reflection of finite order by 
\cite[Proposition 6.4]{KKZ1}. Therefore $A^G$ is AS Gorenstein 
if and only if the $G$-action has trivial homological determinant. 
In \cite[Theorem 3.6]{KKZ3} it is shown that if $H$ is a 
semisimple Hopf algebra acting on an AS regular algebra with 
trivial homological determinant, then $A^H$ is AS Gorenstein; 
it is reasonable to ask whether the converse holds for algebras 
with no ``dual reflection groups".  Hence
Theorem \ref{xxthm0.1} suggests the following question in the 
group coaction setting: 
{\it Let $A$ be a graded noetherian down-up algebra with a finite group
$G$-coaction. If the fixed subring $A^{co\; G}$ is AS Gorenstein, must
the homological determinant of the $G$-coaction be trivial?} 

We conclude the paper with an example (Example \ref{xxex2.1}) that 
provides a negative answer to the above question,  indicating a 
difference between the invariant theory under group actions and the 
invariant theory under group coactions (or more generally, under 
Hopf algebra actions). It would be interesting to develop further 
tools that would determine precisely when,  under a $G$-coaction, 
the fixed subring $A^{co \; G}$ is AS Gorenstein, and, more generally, 
when the  homological determinant of a Hopf action being trivial is a 
necesary condition for $A^H$ to be AS Gorenstein.

\subsection*{Acknowledgments}
J. Chen was partially supported by the National Nature Science 
Foundation of China (Grant No. 11571286)
and the Natural Science Foundation of Fujian Province of China 
(Grant No. 2016J01031). E. Kirkman was partially supported by
grant \#208314 from the Simons Foundation. J.J. Zhang was partially 
supported by the US National Science Foundation (Grant No. DMS 1402863).

\section{Proof of Theorem \ref{xxthm0.1}}
\label{xxsec1}

Some basic definitions can be found in \cite{BB, KKZ5}; for example,
inner faithful is defined in \cite[Definition 2.7]{BB}.
Artin-Schelter regular will be abbreviated by AS regular; for the definition see
\cite[Definition 1.1]{KKZ5}.
We first recall some basic facts about down-up algebras
from \cite{BR, KK, KMP, KKZ4}.

\begin{lemma}
\label{xxlem1.1} Let $A$ be the down-up algebra
$\DD(\alpha,\beta)$ where $\beta\neq 0$.
\begin{enumerate}
\item[(1)]
$A$ is a connected graded noetherian AS regular domain of global
dimension three.
\item[(2)]
The Hilbert series of $A$ is $((1-t)^2(1-t^2))^{-1}$.
\item[(3)]
$\{ u^i(du)^j d^ k\mid i,j,k\geq 0\}$ is a $\Bbbk$-linear basis
of $A$.
\end{enumerate}
\end{lemma}

The following lemma is well-known.

\begin{lemma}
\label{xxlem1.2}
Let $A$ be a noetherian connected graded AS regular algebra of
global dimension three. Then $A$ is generated by either two or three
elements.
\end{lemma}

\begin{proof} This is \cite[Proposition 1.5]{AS} when $A$ is generated
in degree 1 and \cite[Proposition 1.1(i)]{Ste} when $A$ is not generated
in degree 1.
\end{proof}

It is well-known that, for every finite group $G$, a left
$(\Bbbk G)^\ast$-action on an algebra $A$ is equivalent to a right
$G$-coaction on $A$.  Since $\Bbbk$ is algebraically closed of
characteristic zero, if $G$ is a finite abelian group, the Hopf
algebra $\Bbbk G$ is isomorphic to its dual $(\Bbbk G)^\ast$. This
fact implies that a right $G$-coaction on $A$ is equivalent to some
left $G$-action on $A$. We will use these facts freely. The
following lemma is a rigidity result for abelian $G$-coactions.

\begin{lemma}
\label{xxlem1.3}
Let $A$ be a graded noetherian down-up algebra and $G$ be a
nontrivial finite abelian group coacting on $A$. Then the
fixed subring $A^{co\; G}$ is not AS regular.
\end{lemma}

\begin{proof} If $G$ is abelian, $\Bbbk G$ is isomorphic to
$(\Bbbk G)^{\ast}$ as Hopf algebras.
Since $G$ is abelian coacting on $A$, there is a $G$-action on $A$
such that $A^{co\; G}=A^G$. The assertion is a consequence
of \cite[Proposition 6.4]{KKZ1}.
\end{proof}

We consider, first, the case when $\alpha= 0$ and $G$ coacts 
homogeneously on the generators $u$ and $d$.  Note that 
although we show that $\DD_{\beta}$
is rigid with respect to these group coactions, the algebra 
can be graded by many different groups.

\begin{proposition}
\label{xxpro1.4} Let $A$ be the algebra $\DD_{\beta}$ and $G$ be a
nontrivial finite group coacting on $A$ such that $u$ and $d$ are
$G$-homogeneous. Then the fixed subring $A^{co\; G}$ is not AS regular.
\end{proposition}

\begin{proof} By Lemma \ref{xxlem1.3}, we can assume that $G$ is not abelian.
Let $\deg_G u=a$ and $\deg_G d=b$; then $G$ is generated by $\Re:=\{a,b\}$.
Using the relations $u^2 d=\beta du^2$ and $ud^2=\beta d^2 u$, we obtain
that $a^2$ and $b^2$ are in the center of $G$. This implies that
the orders of $a$ and $b$ are even. Let $i$, $j$, $k$ and $l$ be the smallest
positive integers such that $a^i=e$, $b^j=e$, $(ba)^k=e$ and $(ab)^l=e$.
Since $G$ is non-abelian, $i,j,k,l$ are all larger than 1, and both of $i$ and $j$
must be even. Then $x:=u^i$, $y:=d^j$, $z:=(du)^k$ and $t:=(ud)^l$ are
elements in $A^{co\; G}$.

Assume to the contrary that the fixed subring $A^{co\; G}$ is AS regular.
By Lemma \ref{xxlem1.2}, it is generated by at most three elements.
Choosing the generators of $A^{co\; G}$ carefully from lower degree to
higher, and using the fact that every monomial in $u$ and $d$ is
$G$-homogeneous, we can assume that $A^{co\; G}$ is generated by
$$h_s: =u^{m_s} (du)^{n_s} d^{p_s},$$
where $s$ is $1,2,3$ or $1,2$. If $m_s+n_s>0$ for all $s$, then
$y=d^j$ cannot be generated by $\{h_s\}_{s}$, a contradiction. Thus
$m_s+n_s=0$ for some $s$. Similarly, we have $n_s+p_s=0$ for some
$s$. These facts mean that we have
$$h_1=u^i=x \quad {\text{and}}\quad  h_2=d^j=y.$$
Since $x$ and $y$ do not generate $z$, $A^{co\; G}$ is generated
by three elements, namely, by $h_1=u^i$, $h_2=d^j$ and $h_3= u^m (du)^n d^p$.
If $m>0$, then $z=(du)^k$ cannot be generated by $h_1,h_2,h_3$.
Thus $m=0$. By symmetry, $p=0$. This implies that
$$h_3=(du)^n.$$
Since $i$ and $j$ are even, $t=(ud)^l$ cannot be generated by
$h_1,h_2,h_3$. Hence $A^{co\; G}$ is not generated by three (or fewer)
elements. This yields a contradiction by Lemma \ref{xxlem1.2}, and
therefore the fixed subring $A^{co\; G}$ is not AS regular.
\end{proof}

Next we consider an algebra  $\FF$ that is isomorphic to $\DD_{-1}$.
Let $\FF$ be the algebra generated by $x$ and $y$,
and subject to the two relations
\begin{equation}
\label{E1.4.1}\tag{E1.4.1}
x^3=yxy \quad \mbox{ and } \quad y^3=xyx.
\end{equation}

\begin{lemma}
\label{xxlem1.5} Retain the above notation.
\begin{enumerate}
\item[(1)]
$\FF$ is isomorphic to $\DD_{-1}$.
\item[(2)]
$\FF$ is a graded noetherian AS regular domain of global dimension three
with Hilbert series $((1-t)^2(1-t^2))^{-1}$.
\end{enumerate}
\end{lemma}

\begin{proof} (1) Setting $x=u+d$ and $y=u-d$, then the two relations
$x^3=yxy$ and $y^3=xyx$ are equivalent to the two relations
$u^2d=-du^2$ and $ud^2=-d^2u$. The assertion follows.

(2) This follows from the fact that all the assertions hold
for $\DD_{-1}$.
\end{proof}

Next we will apply Bergman's Diamond Lemma \cite{Be} to the algebra
$\FF$. By \cite{Be}, starting
with a set of initial relations, we can obtain a complete set 
of relations that is a
reduction system. Then every monomial (or word) becomes reduction-unique
by using this complete system.

\begin{lemma}
\label{xxlem1.6} Retain the notation as above.
\begin{enumerate}
\item[(1)]
Define an order on monomials by extending $x<y$ lexicographically. 
Then we have a complete set of five relations that is the reduction system
in the sense of \cite[p.180]{Be}.
$$\begin{aligned}
y^3& =xyx,\\
yxy&=x^3,\\
y^2x^3&=xyx^2y,\\
yx^2yx&=x^3 y^2,\\
yx^4&=x^4y.
\end{aligned}
$$
\item[(2)] We also have the other relations:
$$\begin{aligned}
y^4& =x^4,\\
yxyx&=x^4,\\
xyxy&=y^4.
\end{aligned}
$$
\item[(3)]
There is a $\Bbbk$-linear basis consisting of the monomials
of the form
$$x^i (yx^3)^j (yx^2)^{\epsilon} (y^2x^2)^k y^a x^b$$
where $i,j,k\geq 0$, $\epsilon$ is either $0$ or $1$,
and
$$(a,b)=(0,0), (1,0), (1,1), (1,2), (1,3), (2,0), (2,1), (2,2)$$
if $j+\epsilon+k=0$,
$$(a,b)=(1,0), (1,1), (1,2), (1,3), (2,0), (2,1), (2,2)$$
if $j>0$ and $\epsilon+k=0$, and
$$(a,b)=(1,0), (2,0), (2,1), (2,2)$$
if $\epsilon+k>0$.
\end{enumerate}
\end{lemma}

\begin{proof} (1) The assertion follows by direct computation. 
In fact, denote the relation $y^3 =xyx$ by (i), and the relation 
$yxy=x^3$ by (ii). Then, by using (i)+(ii), we have 
$y^3xy=y^2(yxy)=y^2x^3$ and $y^3xy=y^3(xy)=(xyx)(xy)=xyx^2y$. 
Thus we obtain the third relation (denote it by (iii)) in the list.
By using (ii)+(i), we have $yxy^3=(yx)y^3=(yx)xyx=yx^2yx$, 
and $yxy^3=(yxy)y^2=x^3y^2$, and thus the fourth relation 
(denote it by (iv)) in the list holds. Similarly, by using (ii)+(i), 
we obtain the fifth relation (denote it by (v)) in the list.  Then 
considering all the other possible
cases: (i)+(iii), (i)+(iv), (i)+(v),(ii)+(iii), (ii)+(iv) and (ii)+(v), 
there are no new relations.

(2) These assertions follow easily.

(3) Every monomial is of the form
$x^{i_1}y^{j_1}x^{i_2}y^{j_2}\cdots y^{j_{n-1}}x^{i_n}$,
where $n\geq 1$, $i_n\geq 0$, $j_s\geq 1$ and $i_s\geq 1$
for $ 1\leq s<n$. By the first relation $j_s$ can  be only 
1 or 2, and by the fifth relation in part (1), for $s>1$, 
$i_s$ can  be only 1, 2, or 3.

Take the last term $n$ into consideration. If $n=1$, we only have $x^{i_1}$. If $n=2$, we 
have only
$$x^i y, \quad x^i y^2, \quad x^i yx, \quad x^i yx^2,
\quad x^i yx^3, \quad x^i y^2 x, \quad x^i y^2 x^2$$
or
$$x^i y^{a} x^b$$
where $i\geq 0$ and 
\begin{equation}
\notag
(a,b)=(1,0), (1,1), (1,2), (1,3), (2,0), (2,1), (2,2).
\end{equation}
For $n\geq 3$, note that in the middle of the monomial,
$y^{j_{s-1}} x^{i_s}$ must be $yx^2$, or $yx^3$, or $y^2 x^2$.
Further, $yx^2 yx^3$, $y^2x^2 yx^3$, $y^2x^2 y x^2$, $y x^2 y x^2$
cannot appear in the middle of the monomial. Therefore we have
$$x^i (yx^3)^j (yx^2)^{\epsilon} (y^2x^2)^k y^a x^b$$
where $i,j,k\geq 0$, $\epsilon$ is either $0$ or $1$,
and
$$(a,b)=(0,0), (1,0), (1,1), (1,2), (1,3), (2,0), (2,1), (2,2)$$
if $j+\epsilon+k=0$,
$$(a,b)=(1,0), (1,1), (1,2), (1,3), (2,0), (2,1), (2,2)$$
if $j>0$ and $\epsilon+k=0$, and
$$(a,b)=(1,0), (2,0), (2,1), (2,2)$$
if $\epsilon+k>0$.
By Bergman's Diamond Lemma \cite{Be}, all monomials in part (3) form a
$\Bbbk$-linear basis of the algebra.
\end{proof}

Let $f$ be a monomial $x_1 x_2 x_3\cdots x_n$ in $\FF$ where $x_i$
is either $x$ or $y$. A {\it left subword} of $f$ is a monomial
of the form $x_1 x_2\cdots x_j$ for $j\leq n$, a {\it subword}
of $f$ is a monomial of the form $x_i\cdots x_j$ for some
$1\leq i\leq j\leq n$. Due to non-trivial relations in $\FF$,
if $ab=f$ for three monomials $a,b,f$, it is not necessarily true
that $a$ is a left subword of $f$.  The following lemma says that
in some special cases, $a$ must be a left subword of $f$.

\begin{lemma}
\label{xxlem1.7} Let $f$ be a subword of $(y^2 x^2)^s=yyxxyyxx\cdots yyxx$
for some $s\geq 1$. If $f=ab$ for some monomials $a,b$ in $\FF$, then
$a$ is a left subword of $f$.
\end{lemma}

\begin{proof} By changing $s$ to a larger number and adding more letters
to $f$ and $b$ from the right, we may assume that the degree of $f$ is at least
4.  We prove the assertion by induction on the degree of $a$. Nothing
needs to be proved if $a$ has degree 0. Now suppose $\deg a>0$. There are
four different cases for $f$:
$$f=yyxx\cdots, \quad yxxy\cdots, \quad xxyy\cdots, \quad  xyyx\cdots.$$
By Lemma \ref{xxlem1.6}(3), each $f$ is reduced. Suppose $f=yf'$ is in the first
two cases. If $a=ya'$, then, after canceling out $y$, we have $f'=a'b$, and the assertion
follows from the induction. If $a=xa'$, then the reduced form of $ab$ is
less than $f$ in the order used in the  Diamond Lemma, but this is impossible, and so we are done
in this case.  Suppose next that $f=xf'$ is in one of the last two cases. If $a=xa'$,
then, after canceling out $x$, we have $f'=a'b$, and the assertion
follows from the induction. The remaining case is $a=ya'$, and we need to consider the following two
separate cases for $f$.

If $f=xxyy\cdots$ and $a=ya'$, then $yyf$ is a subword of $(y^2x^2)^{s+1}$ and
$y^2ab=y^3 a'b=xyx a'b$, which is less than $yyf$ in the order used in the Diamond
Lemma, but this is impossible.
If $f=xyyx\cdots$ and $a=ya'$, then $yxf=yxxyyx\cdots$ and $yxya'b=
x^3 a'b$, which is less than $yxxyyx\cdots=yxf$ in the order used in the Diamond
Lemma, but this is impossible.

Combining the above assertions, the induction shows that $a$ is a left subword of 
the word $f$.
\end{proof}

\begin{proposition}
\label{xxpro1.8}
Let $A$ be the algebra $\FF$ and $G$ be a nontrivial finite group
coacting on $A$ such that $x$ and $y$ are $G$-homogeneous. Then
the fixed subring $A^{co\; G}$ is not AS regular.
\end{proposition}

\begin{proof} By Lemma \ref{xxlem1.3}, we need to
consider only the case when $G$ is not abelian. Assume to the contrary that
$\FF^{co\; G}$ is AS regular. By \cite[Lemma 3.3(2)]{KKZ5},
$\FF^{co\; G}$ has global dimension three.
Note that $\FF$ is a semigroup
algebra $\Bbbk T$ for the semigroup
$$T=\langle x,y\mid x^3=yxy, y^3=xyx\rangle$$
and $G$ is a finite factor group $T/N$ for some normal
subsemigroup $N$. We have
\begin{enumerate}
\item[(1)]
The fixed subring $\FF^{co\; G}$ is the semigroup ring $\Bbbk N$.
\item[(2)]
$\FF^{co\; G}$ is minimally generated by a finite subset
$S\subset N$.
\item[(3)]
Every monomial in $\FF^{co \; G}$ is a product of elements in $S$.
\end{enumerate}
Note that we have identified a monomial in $\FF$ with an element in
$T$. Let $g_1$ be the image of $x$ in $G$ and $g_2$ be the image
of $y$ in $G$. Then there is an $s>1$ such that
$(g_2g_2 g_1g_1)^s\in N$, or equivalently,  $(g_2g_2 g_1g_1)^s=e$
in $G$. Then $f_1=(yyxx)^s\in \FF^{co\; G}$. Similarly we have
three other monomials in $\FF^{co\; G}$:
$$\begin{aligned}
f_2&= yxx(yyxx)^{s-1} y,\\
f_3&= xx(yyxx)^{s-1} yy,\\
f_4&= x(yyxx)^{s-1} yyx.
\end{aligned}
$$
Then $f_1,f_2,f_3,f_4$ are four elements in $\FF^{co\; G}$.
Since $f_1$ is in $\FF^{co\; G}$, $f_1=a_1 b_1$ where $a_1$
is in the set $S$. Similarly, we have $a_i \in S$ such that
$f_i=a_i b_i$ for $i=2,3,4$. By Lemma \ref{xxlem1.7},
$a_i$ is a left subword of $f_i$. By Lemma \ref{xxlem1.6}(3),
as left subwords of $f_1,f_2,f_3,f_4$ respectively, $a_1,a_2,a_3,a_4$,
are reduced and linearly independent. Therefore the order of
$S$ is at least 4, which contradicts Lemma \ref{xxlem1.2}. Therefore
$\FF^{co\; G}$ is not AS regular.
\end{proof}

We consider another algebra $\HH$ that is generated by
$x$ and $y$ subject to the relations
$$\begin{aligned}
x^2 y+yx^2 -2y^3&=0,\\
-2x^3+xy^2+y^2 x&=0.
\end{aligned}
$$

The following lemma is similar to Lemma \ref{xxlem1.5}.

\begin{lemma}
\label{xxlem1.9} Retain the above notation.
\begin{enumerate}
\item[(1)]
$\HH$ is isomorphic to $\DD(-2,-1)$.
\item[(2)]
$\HH$ is a graded noetherian AS regular domain of global dimension three
with Hilbert series $((1-t)^2(1-t^2))^{-1}$.
\item[(3)]
For each $n>0$, elements $(xy)^n$ and $(yx)^n$ are linearly independent.
\item[(4)]
For each $n\geq 0$, elements $y(xy)^n$ and $x(yx)^n$ are linearly independent.
\end{enumerate}
\end{lemma}

\begin{proof} (1) Set $u=x-y$ and $d=x+y$, then the two relations
$u^2d=-2udu-du^2$ and $ud^2=-2dud-d^2u$ are equivalent to the two relations
$x^2 y+yx^2 -2y^3=0$ and
$-2x^3+xy^2+y^2 x=0$.

(2) Known for $\DD(\alpha, \beta)$ for all $\beta\neq 0$.

(3) For any $d \geq 2$, let $D_{2d}$ be the dihedral group of order $2d$ 
generated by $a$ and $b$ subject to the relations $a^2=b^2=(ab)^d=e$
where $e$ is the unit of $D_{2d}$.
Consider the $G:=D_{2d}$-coaction on $A$ obtained by setting
$\deg_G x=a$ and $\deg_G y=b$.  It is clear that $A$ is $G$-graded.  Choosing $d\gg n$, then
$$\deg_G((xy)^n)=(ab)^n\neq (ba)^n =\deg_G ((yx)^n),$$
which implies the assertion.

(4) The proof is similar to the proof of part (3), and so it is omitted.
\end{proof}

\begin{proposition}
\label{xxpro1.10}
Let $A$ be the algebra $\HH$ and $G$ be a nontrivial finite group
coacting on $A$ such that $x$ and $y$ are $G$-homogeneous. Then
the fixed subring $A^{co\; G}$ is not AS regular.
\end{proposition}

\begin{proof} By Lemma \ref{xxlem1.3}, we may assume that
$G$ is non-abelian. Suppose to the contrary that $A^{co\; G}$ is
AS regular. Let $g_1:= \deg_G x$ and $g_2:=\deg_G y$.  The 
$G$-grading forces $g_1^2=g_2^2$.  Let $a:=g_1^2=g_2^2$. If 
$a\neq e$, then both $x^2$ and $y^2$
are in the same $G$-graded component $A_{a}$. By \cite[Theorem 0.3(1)]{KKZ5},
$A_{a}$ is free of rank 1 over
$A^{co\; G}$. Thus $A_a=x^2 A^{co\; G}$ and
$A_a=y^2 A^{co\; G}$. This contradicts the fact that $x^2$ and $y^2$ are
linearly independent. Therefore $g_1^2=g_2^2=e$. As a consequence,
$G$ is isomorphic to the dihedral group $D_{2n}$ of order $2n$,
for some $n \geq 2$. When $n$ is odd the unique largest length element 
of $G$ with respect to the Coxeter generating set $\{g_1, g_2\}$ is
$m=(g_1 g_2)^{(n-1)/2} g_1=(g_2g_1)^{(n-1)/2} g_2$,
while when $n$ is even it
is $m=(g_1 g_2)^{n/2} =(g_2g_1)^{n/2}$, which is central.
By Lemma \ref{xxlem1.9}(3,4),
when $n$ is odd, $(xy)^{(n-1)/2} x$ and $(yx)^{(n-1)/2}y$
are linearly independent elements of degree $n$ in the $G$-graded 
component $A_m$, and when $n$ is even, $(xy)^{n/2}$ and $(yx)^{n/2}$
are linearly independent elements of degree $n$ in $A_m$. But the smallest
degree of elements in $A_m$ is $n$. This yields a
contradiction by \cite[Theorem 0.3(4)]{KKZ5}.
\end{proof}

For the rest of this section we assume that $G$ is a finite
group coacting on the down-up algebra $\DD(\alpha,\beta)$.

\begin{lemma}
\label{xxlem1.11} 
Let $A$ be the connected graded algebra $\DD(\alpha,\beta)$
with $G$-coaction. Let $x_1$ and $x_2$ be two linearly 
independent $G$-homogeneous elements in $A_1$. Suppose there 
are two nontrivial relations that hold in $A$:
$$f_1:=\sum_{i,j,k=1}^2 c_{i,j,k} x_i x_j x_k=0,$$
and
$$f_2:=\sum_{i,j,k=1}^2 e_{i,j,k} x_i x_j x_k=0$$
such that $c_{i,j,k}\neq 0$ and $e_{i,j,k}=0$
for some $(i,j,k)$. Then,  for
all $(i,j,k)$ with $e_{i,j,k}\neq 0$, the 
$x_ix_jx_k$ have the same $G$-degree.
\end{lemma}

\begin{proof}
Since the monomial $x_ix_jx_k$ does not appear in the relation 
$f_2$ with nonzero coefficient, $f_1$ and $f_2$ are linearly 
independent.  If $f_2$ contains two monomials with nonzero 
coefficients and different $G$-degrees, then $f_2$ must be a sum 
of  $G$-homogeneous pieces $g_1, \dots, g_n$ for $n \geq 2$, with 
each $G$-homogeneous piece a relation in $A$ of degree 3.  But 
then $f_1, g_1$, and $g_2$ are three linearly independent relations 
in $A$ of degree 3, which is a contradiction.
\end{proof}

\begin{proposition}
\label{xxpro1.12}
Suppose $G$ is a finite non-cyclic group coacting on
$A:=\DD(\alpha,\beta)$ homogeneously and inner faithfully.
Then one of the following occurs.
\begin{enumerate}
\item[(1)]
$\alpha=0$ and $u$ and $d$ are $G$-homogeneous after
a change of variables.
\item[(2)]
$A$ is isomorphic to $\FF$ and using the generators of $\FF$,
both $x$ and $y$ are $G$-homogeneous.
\item[(3)]
$A$ is isomorphic to $\HH$ and using the generators of $\HH$,
both $x$ and $y$ are $G$-homogeneous.
\item[(4)]
$G$ is abelian and there are linearly independent elements
$x$ and $y$ of $\DD(\alpha, -1)$ of degree one
such that
$$\begin{aligned}
\alpha x^2 y+(-2-\alpha) xyx+ \alpha yx^2 +(2-\alpha) y^3&=0,\\
(2-\alpha) x^3+ \alpha xy^2+(-2-\alpha) yxy+ \alpha y^2 x&=0
\end{aligned}
$$
and  $x$ and $y$ are $G$-homogeneous.
\item[(5)]
$G$ is abelian and $u$ and $d$ are $G$-homogeneous after a change
of variables.
\end{enumerate}
\end{proposition}

\begin{proof} Each of the five parts listed in
 (1-5) can occur. Part (5) could occur most often, so,
for the rest of the proof, we implicitly assume that we are
not in the situation of {\bf part (5)}.

Write $A_1=\Bbbk x+ \Bbbk y$ where $x,y$ are $G$-homogeneous.  
Then $g_1:=\deg_G x$ and $g_2:=\deg_G y$ generate $G$. Since 
$G$ is not cyclic, we have
$$e \neq g_1\neq g_2\neq e.$$

{\bf Case 1 $\alpha = 0, \beta=1$:} First, we assume that $A=\DD_1$.
Then, for any two linearly independent elements $x,y$ of $A$
of degree 1, one can check that $x^2$ and $y^2$ are central.
Therefore we can assume that $x$ and $y$ are $G$-homogeneous,
by the second paragraph. After changing $\{u,d\}$ to
$\{x,y\}$, we can assume that $u$ and $d$ are $G$-homogeneous.
Thus {\bf part (1)} holds for  $A=\DD_1$.

{\bf Case 2 $\alpha = 0, \beta \neq 1$:} Secondly, 
we assume that $A=\DD_{\beta}$ where $\beta\neq 1$.
As noted in the second paragraph, there are two elements $x$ and $y$
in degree 1 with different $G$-grades. We consider two cases.

{\bf Case 2a:}  $x=cu$ and $y=au+b d$ for some $a,b,c\in \Bbbk$. Since $x$ and
$y$ are linearly independent, $bc\neq 0$. If $a=0$, then we can choose
$x=u$ and $y=d$ after a change of variables, and the assertion in
{\bf part (1)} follows. Now we assume $abc\neq 0$. Up to another change
of variables, we have $x=u$ and $y=u+d$, or equivalently, $u=x$
and $d=y-x$. Then the two relations \eqref{E0.0.1} and \eqref{E0.0.2} become
\begin{align}
\label{E1.12.1}\tag{E1.12.1}
x^2 y-\beta y x^2+(\beta-1) x^3&=0,\\
\label{E1.12.2}\tag{E1.12.2}
xy^2 -\beta y^2 x+(1-\beta) x^3&=x^2 y+(1-\beta)xyx -\beta yx^2.
\end{align}
Combining these two relations, one obtains that
\begin{equation}
\label{E1.12.3}\tag{E1.12.3}
xy^2 -\beta y^2 x+(\beta-1)xyx=0.
\end{equation}
Then \eqref{E1.12.1} must be
$G$-homogeneous by Lemma \ref{xxlem1.11}. As a
consequence, $x$ and $y$ have the same $G$-grade,
which contradicts the fact that $G$ is not cyclic.

{\bf Case 2b:} Up to a change of variables, the remaining
case is when $u=x-y$ and $d=x-ay$ where $a\neq 0,1$.
The two relations \eqref{E0.0.1} and \eqref{E0.0.2} become
$$(1-\beta)x^3+(-a+\beta)x^2y+(-1+\beta)xyx
+(-1+a\beta)yx^2$$
$$+(a-\beta)xy^2
+(a-a\beta) yxy +(1-a\beta) y^2 x+(-a+a\beta)y^3=0$$
and
$$(1-\beta)x^3+(-a+\beta)x^2y+(-a+a\beta)xyx
+(-1+a\beta) yx^2$$
$$+(a^2-a\beta)xy^2+(a-a\beta)yxy+(a-a^2\beta)y^2x
+(-a^2+a^2\beta)y^3=0.$$
Since $a \neq 1$, by linear combination, we have
$$(1-\beta)x^3+(-a+\beta)x^2y+(-1+a\beta)yx^2+a(1-\beta) yxy=0$$
and
$$(1-\beta)xyx+(-a+\beta)xy^2+(-1+a\beta)y^2x
+a(1-\beta)y^3=0.$$
Note that $1-\beta\neq 0$. Since $G$ is not cyclic,
$-a+\beta=0$ and $-1+a\beta=0$ by Lemma \ref{xxlem1.11}.
Thus $\beta=a=-1$, and we have the relations $$x^3-yxy=0 \mbox{ and } y^3-xyx=0$$
which is {\bf part (2)}.

{\bf Case 3 $\alpha \beta \neq 0$:} Thirdly, we assume that $A=\DD(\alpha,\beta)$ where
$\alpha \beta\neq 0$. As before we  need
to consider two cases by Lemma \ref{xxlem1.11}.

{\bf Case 3a}: Let $x=u$ and $y=u+d$, or equivalently
$u=x$ and $d=y-x$. The relations \eqref{E0.0.1}
and \eqref{E0.0.2} become
$$x^2y -\alpha xyx -\beta yx^2+(-1+\alpha+\beta) x^3=0$$
and
$$(1-\alpha-\beta)x^3+(-1+\alpha)x^2y+(-1+\beta)xyx$$
$$+(\alpha+\beta) yx^2+ xy^2+(-\alpha) yxy+(-\beta)y^2 x=0.$$
By adding these two relations, we obtain that
$$\alpha x^2y +(-1-\alpha+\beta) xyx+\alpha yx^2
+ xy^2-\alpha yxy-\beta y^2 x=0.$$
If $-1+\alpha+\beta\neq 0$, this forces  $\deg_G(x^2y)=\deg_G(xy^2)$.
This implies that $\deg_G(x)=\deg_G(y)$ and $G$ is cyclic,
a contradiction. If $-1+\alpha+\beta=0$, we still have
$\deg_G(x^2y)=\deg_G(xyx)$, which implies that $G$ is abelian.
Further, by Lemma \ref{xxlem1.11}, up to a common scalar,
the relation 
$$x^2y -\alpha xyx -\beta yx^2=0$$ 
must coincide with the relation 
$$\alpha x^2y +(-1-\alpha+\beta) xyx+\alpha yx^2=0.$$
As a consequence, $(\alpha, \beta)=(2,-1)$, and we have (a special case of) 
{\bf part (4)}.

{\bf Case 3b:} The remaining case is when $u=x-y$ and $d=x-ay$ where
$a\neq 0,1$. Using these generators to expand the relations \eqref{E0.0.1}
and \eqref{E0.0.2} we have
$$\begin{aligned}
(1-\alpha-\beta)x^3&+(-a+\alpha+\beta) x^2y
+(-1+a\alpha+\beta) xyx+(-1+\alpha+a\beta) yx^2\\
+(a-a\alpha-\beta) &xy^2+(a-\alpha-a\beta) yxy
+(1-a\alpha-a\beta)y^2 x+a(-1+\alpha+\beta) y^3=0\\
\end{aligned}$$
{and}\\
$$\begin{aligned}
(1-\alpha-\beta)x^3&+(-a+a\alpha+\beta)x^2y
+(-a+\alpha+a\beta) xyx+(-1+a\alpha+a\beta)yx^2+\\
a(a-\alpha-\beta) &xy^2+a(1-a\alpha-\beta) yxy
+a(1-\alpha-a\beta)y^2x+a^2(-1+\alpha+\beta)y^3
=0.
\end{aligned}
$$
If $1-\alpha-\beta=0$, then the  two relations above
become
$$\begin{aligned}
(1-a) x^2y&
+\alpha(a-1) xyx+\beta(a-1) yx^2\\
&+\beta(a-1) xy^2+\alpha(a-1) yxy
+(1-a)y^2 x=0\\
\end{aligned}$$
and
$$\begin{aligned}
\beta(1-a)x^2y&
+\alpha(1-a) xyx+(a-1)yx^2\\
&+a(a-1) xy^2+a\alpha(1-a) yxy
+a\beta(1-a)y^2x   =0.  
\end{aligned}
$$
Since $a\neq 1$, after dividing by $1-a$, we have
$$x^2y-\alpha xyx-\beta yx^2
-\beta xy^2-\alpha yxy
+y^2 x=0$$
and
$$\beta x^2y +\alpha xyx-yx^2
-a xy^2+a\alpha yxy
+a\beta y^2x=0
$$
which are equivalent to
$$\begin{aligned}
(1+\beta)x^2y+(-1-\beta)yx^2
+(-\beta-a)xy^2+\alpha (a-1) yxy
+(1+a\beta)y^2 x&=0\\
\end{aligned}$$
and
$$\begin{aligned}
(a+\beta) x^2y
+\alpha(1-a)xyx+(-a\beta-1)yx^2
+(-a\beta -a) xy^2+(a+a\beta) y^2x&=0.
\end{aligned}
$$
If $\beta\neq -1$, by Lemma \ref{xxlem1.11}, we have
$\deg_G(xyx)=\deg_G(y^2x)$, which implies that
$G$ is cyclic, a contradiction. If $\beta=-1$
(and then $\alpha=2$),
then the above two relations become
$$\begin{aligned}
xy^2-2 yxy +y^2 x&=0,\\
x^2y-2 xyx+ yx^2&=0,
\end{aligned}
$$
so by Lemma \ref{xxlem1.11} $G$ is abelian. Therefore
we have {\bf part (4)}.

If $1-\alpha-\beta\neq 0$, the two relations given at the
beginning of Case 3b are equivalent to
$$\begin{aligned}
\alpha(1-a)x^2y+&(a-1)(1+\alpha-\beta)xyx+\alpha (1-a) yx^2+(1-a)(a-\beta) xy^2\\
+\alpha(a^2-1) & yxy+(1-a)(1-a\beta) y^2x
+a(1-a)(-1+\alpha+\beta) y^3=0,\\
\end{aligned}$$
and
$$\begin{aligned}
(1-a)(1-\alpha-\beta)&x^3+(a-1)(a-\beta)x^2 y +\alpha(1-a^2) xyx+
(1-a)(-1+a\beta) yx^2\\
+a\alpha(a-1) &xy^2+ a(1-a) (1+\alpha-\beta) yxy+ a\alpha (a-1) y^2x=0 .
\end{aligned}
$$
Since $a\neq 1$, we can simplify them to obtain the following two relations
$$\begin{aligned}
\alpha x^2y-(1+\alpha-\beta)xyx+&\alpha yx^2+(a-\beta) xy^2-\alpha(a+1)  yxy\\
 &+(1-a\beta) y^2x+a(-1+\alpha+\beta)y^3=0,
\end{aligned}
$$
and
$$\begin{aligned}
 (1-\alpha-\beta)x^3-(a-\beta)x^2 y +&\alpha(1+a) xyx+(-1+a\beta) yx^2\\
&-a\alpha xy^2
+ a (1+\alpha-\beta) yxy- a\alpha  y^2x=0 .
\end{aligned}
$$

Suppose $a\neq -1$. Since $\alpha (a-1)\neq 0$, the coefficients
of $yxy$ and $y^3$ are nonzero. By Lemma \ref{xxlem1.11},
we have $\deg_G(yxy)=\deg_G(y^3)$. This forces $G$ to be cyclic,
a contradiction. If $a=-1$ and $\beta\neq -1$, a similar argument
leads to a contradiction. If $a=-1=\beta$, we obtain two
relations
$$\begin{aligned}
\alpha x^2 y+(-2-\alpha) xyx+ \alpha yx^2 +(2-\alpha) y^3&=0,\\
(2-\alpha) x^3+ \alpha xy^2+(-2-\alpha) yxy+ \alpha y^2 x&=0.
\end{aligned}
$$
In addition, if $\alpha\neq -2$, then the coefficients of $x^2y$ and $xyx$ are nonzero. 
By Lemma
\ref{xxlem1.11}, $\deg(x^2y)=\deg(xyx)$ which implies that $G$ is abelian, and so we have {\bf part (4)};
 if $\alpha=-2$, the two relations become
$$\begin{aligned}
 x^2 y+ yx^2 -2y^3&=0,\\
-2 x^3+ xy^2+ y^2 x&=0,
\end{aligned}
$$
which is {\bf part (3)}. This completes the proof.
\end{proof}

\begin{proof}[Proof of Theorem \ref{xxthm0.1}]
If $G$ is abelian, then the assertion follows from
Lemma \ref{xxlem1.3}. We now assume that $G$ is non-abelian.
By Proposition \ref{xxpro1.12} there are three cases
(parts (1), (2) and (3) occur). The first case follows from
Proposition \ref{xxpro1.4}, the second case follows
from Proposition \ref{xxpro1.8}, and the third case follows
from Proposition \ref{xxpro1.10}.
\end{proof}

\section{An Example}
\label{xxsec2}

In this short section we provide an example that negatively 
answers the question  mentioned at the end of the introduction.

\begin{example}
\label{xxex2.1}
Consider the algebra $\DD:=\DD(0,1)$, and
let $G$ be the dihedral group 
$$D_{8}=\{e, \rho, \rho^2, \rho^3, r, \rho r, \rho^2 r, \rho^3 r\},$$
where $\rho$ is a rotation of order $4$ and $r$ is a reflection.
Let $g_1=r$ and $g_2=\rho$. Then $\{g_1,g_2\}$ generate the group
$G$. Define the $G$-coaction on $\DD$ by setting $\deg_G u=\rho$
and $\deg_G d=r$. By a computation similar to \cite[Example 7.1]{KKZ3},
one sees that the homological co-determinant of the $G$-coaction is 
$g_1^2 g_2^2=\rho^2$, which is not trivial. In other words, 
the $(\Bbbk G)^{\circ}$-action on $\DD$ has nontrivial 
homological determinant. On the other hand, by Lemma \ref{xxlem2.2}
below, the fixed subring $\DD^{co \; G}$ is 
$\Bbbk [(du)^2, (ud)^2,d^4, u^2]$ which is isomorphic to 
$\Bbbk [x,y,z,t]/(xy-zt^2)$.  
As a consequence, $\DD^{co\;  G}$ is a commutative complete 
intersection (and hence AS Gorenstein), but not AS regular. 
\end{example}

\begin{lemma}
\label{xxlem2.2}
Retain the notation as the above example. Let $\DD_{g}$ be the
$g$-component of $\DD$ for all $g\in D_{8}$.
\begin{enumerate}
\item[(1)]
$\DD_{e}=\DD^{co\; G}=\Bbbk [(du)^2, (ud)^2,d^4, u^2]$.
\item[(2)]
$\DD_{\rho}=u \DD^{co\;  G} \cong \DD^{co\;  G} (-1)$. 
\item[(3)]
$\DD_{\rho^2}=u^2 \DD^{co\;  G} \cong \DD^{co\;  G} (-2)$. 
\item[(4)] 
$\DD_{\rho^3}=u^3 \DD^{co\;  G}+dud \DD^{co\;  G}$.
\item[(5)]
$\DD_{r}=d \DD^{co\;  G}+udu \DD^{co\;  G}$.
\item[(6)]
$\DD_{\rho r}=ud \DD^{co\;  G}+u^2 du \DD^{co\;  G}\cong u \DD_{r}$.
\item[(7)]
$\DD_{\rho^2 r}=u^2 d \DD^{co\;  G}+u^3 du \DD^{co\;  G} \cong u^2 \DD_{r}$.
\item[(8)]
$\DD_{\rho^3 r}=u^3 d\DD^{co\;  G}+du \DD^{co\;  G}$.
\end{enumerate}
\end{lemma}

\begin{proof}
By Lemma \ref{xxlem1.1}(3),
every $G$-homogeneous element in $\DD$ is a linear combination 
of monomials of the form $u^i (du)^j d^k$ for some $i,j,k\geq 0$.  
Note that 
$$\deg_G d^k=\begin{cases} r & {\text{ $k$ is odd}}\\
1& {\text{ $k$ is even}}\end{cases}, \quad {\text{and}}\quad
\deg_G (du)^j=\begin{cases} r\rho & {\text{ $j$ is odd}}\\
1& {\text{ $j$ is even}}\end{cases}.$$
This implies that
$$\deg_G (du)^j d^k=\begin{cases} \rho^3 & {\text{ $j$ and $k$ are odd}}\\
r& {\text{ $j$ is even and $k$ is odd}}\\
r\rho &{\text{ $j$ is odd and $k$ is even}}\\
1 & {\text{ $j$ and $k$ are  even.}}
\end{cases}
$$
Hence
\begin{equation}
\label{E2.2.1}\tag{E2.2.1}
\deg_G u^i (du)^j d^k=\begin{cases} \rho^{i+3} & {\text{ $j$ and $k$ are odd}}\\
\rho^i r& {\text{ $j$ is even and $k$ is odd}}\\
\rho^i  r\rho=\rho^{i-1} r &{\text{ $j$ is odd and $k$ is even}}\\
\rho^i & {\text{ $j$ and $k$ are  even.}}
\end{cases}
\end{equation}

(1) It is clear that elements $(du)^2, (ud)^2,d^4, u^2$ are in
the fixed subring. If $u^i (du)^j d^k$ is in $\DD^{co\; G}$, then 
formula \eqref{E2.2.1} shows that
this could happen only when $j$ and $k$ are odd and $i\equiv 1 \mod 4$, 
or when $j$ and $k$ are  even and $i\equiv 0 \mod 4$. The assertion
follows.

(2) If $u^i (du)^j d^k$ is in $\DD_{\rho}$, then formula \eqref{E2.2.1} 
shows that
this could happen only when $j$ and $k$ are odd and $i+3\equiv 1 \mod 4$, 
or when $j$ and $k$ are  even and $i\equiv 1 \mod 4$. In both cases, $i\geq 1$.
Thus $u^i (du)^j d^k\in u\DD$, and, as a consequence, $u^i (du)^j d^k\in u\DD^{co\;  G}$.

(3) The proof is similar to the proof of (1,2).

(4) Continue the computation in the proof of parts (1,2), if 
$\deg_G u^i (du)^j d^k=\rho^3$, then this could  happen only when 
$j$ and $k$ are odd and $i+3\equiv 3 \mod 4$, 
or when $j$ and $k$ are  even and $i\equiv 3 \mod 4$. In the first case,
$i$ could be 0 and $u^i (du)^j d^k\in dud \DD$; and in the second case,
$i$ could be 3 and $u^i (du)^j d^k\in u^3 \DD$. This implies that
$u^i (du)^j d^k\in u^3 \DD^{co\;  G}+dud \DD^{co\;  G}$.

(5-8) The proofs are similar to the proof of (4) and omitted.
\end{proof}

\end{document}